OpenStudy (anonymous):

If spine bone density is normally distributed in young women with a mean of 1.0 g/cm2 and a standard deviation of .10 g/cm2, then how many young women out of 100 would you expect to have bone densities greater than 1.15 g/cm2?

OpenStudy (anonymous):

\[P(X>1.15)=P\left(\frac{X-1.0}{0.10}>\frac{1.15-1.0}{0.10}\right)=P(Z>1.5)\]